# Algebraic Topology Principles behind Topological Quantum Error Correction

### v1

@@ IV.2 Implementing 3-Torus Code: Encoding on Edges
Thus, the above vertex and face stabilizers are well-defined. And similar to the original toric code [21], a $\sigma_{X}$ error happening on any qubit will cause 2 mismatched $A_{i}^{1}$ stabilizers, where the error correction process can be done by various matching algorithms, such as Maximum Likelihood matching or Minimum Weight Perfect Matching [37]. This process is similar to the original toric code, albeit it is more complicated, arising from the 3-dimensional lattice structure. And a $\sigma_{Z}$ error happening on any qubit will cause 4 mismatched $B_{j}^{1}$ stabilizers, because each edge is adjacent to 4 different faces, additional $\sigma_{Z}$ errors will create the mistakes in pairs, and we can do the matching accordingly, which is again more complicated.
Figure 7 demonstrates an example of the failures of stabilizer checks when several physical qubits are subjected to $\sigma_{X}$ or $\sigma_{Z}$ errors. Figure 7(a) shows 3 qubits suffer from $\sigma_{X}$ error labelled blue. Then, there will be 2 failures in the vertex stabilizer $A^{1}_{\alpha}$ checks, where the corresponding vertices are labelled red. Figure 7(b) shows that 2 qubits suffer from $\sigma_{Z}$ error labelled blue, and in this case, there will be 6 failures in the face stabilizers, $B^{1}_{\alpha}$, where the corresponding faces are labelled red. Notice that for both cases, the stabilizer failures always appear in an even number of sites, thus, the error correction process can be theoretically done via different matching functions, as discussed above [37, 11].
Figure 6 illustrates the logical $\sigma_{X}$ and $\sigma_{Z}$ graphically. The blue lines in Figure 6 forms a non-contractible loop $\gamma^{1}\in C_{1}(T^{3};\mathbb{Z}_{2}),\partial\gamma^{1}=0$ and $[\gamma^{1}]\neq 0$. The red lines forms a non-compressible 2-torus $\gamma^{2}\in C_{2}(T^{3};\mathbb{Z}_{2}),\partial\gamma^{2}=0$ and $[\gamma^{2}]\neq 0$. And they intersect at exactly 1 site, which is the edge drawn in green. These properties allow us to represent a logical $\sigma_{X}$ on the blue qubits and a logical $\sigma_{Z}$ on the red qubits.
@@ IV.3 Implementing 3-Torus Code: Encoding on Faces
Figure 8: The stabilizer checks defined on the 3-torus code for 2-dimensional cells. For each edge, we check the $\sigma_{Z}$ of 4 qubits reside on adjacent faces, the $A_{i}$ stabilizers, drawn in blue. For each cube, we check the $\sigma_{X}$ of 6 qubits reside on the boundary faces, the $B_{j}$ stabilizers, drawn in red.
@@
We have shown above that the 3-torus code can encode 3 qubits on the 1-cells. Here, we demonstrate that the 3-torus code can encode 3 extra qubits on the 2-cells, or faces. Remark that the 3 qubits encoded on the faces are independent of the 3 qubits that are on the edges.
@@
The edge stabilizer $A^{2}_{i}$ is defined to check the $\sigma_{Z}$ of the $4$ qubits on the faces adjacent to a given edge, an example is drawn in blue in Figure 8. And the cube stabilizer $B^{2}_{j}$ is defined to check the $\sigma_{X}$ of the $6$ qubits on the boundary faces of a given cube, an example is drawn in red in Figure 8. $A_{i}^{2}$ and $B_{j}^{2}$ commute for all $i,j$, because they either do not have any qubits in common or have 2 qubits in common. Not surprisingly,
@@
Here, we consider the case of qubits encoded on edges. In a general $u\times v\times w$ lattice for $u,v,w\in\mathbb{Z}^{+}$, there are in total $4uvw$ edges in the lattice. Hence, there will be $4uvw$ physical qubits and three logical qubits can be encoded, where the $\sigma_{Z}$ distance of the 3 qubits equals $u,v$, and $w$ respectively. Furthermore, the $\sigma_{X}$ distance of the 3 qubits equals $vw,wu$, and $uv$ respectively, where Figure 6 has demonstrated the idea. In the special case of a cubic lattice with dimensions $N\times N\times N$, there are $4N^{3}$ edges, the distance for logical $\sigma_{Z}$ is $N$, and the distance for logical $\sigma_{X}$ is $N^{2}$. Hence, the code is a $[[4N^{3},3,N]]$.

### v2

@@ IV.2 Implementing 3-Torus Code: Encoding on Edges
Thus, the above vertex and face stabilizers are well-defined. And similar to the original toric code [21], a $\sigma_{X}$ error happening on any qubit will cause 2 mismatched $A_{i}^{1}$ stabilizers, where the error correction process can be done by various matching algorithms, such as Maximum Likelihood matching or Minimum Weight Perfect Matching [40]. This process is similar to the original toric code, albeit it is more complicated, arising from the 3-dimensional lattice structure. And a $\sigma_{Z}$ error happening on any qubit will cause 4 mismatched $B_{j}^{1}$ stabilizers, because each edge is adjacent to 4 different faces. Additional $\sigma_{Z}$ errors will create mistakes in pairs, and we can do the matching accordingly, which is again more complicated.
Figure 8 demonstrates an example of the failures of stabilizer checks when several physical qubits are subjected to $\sigma_{X}$ or $\sigma_{Z}$ errors. Figure 8(a) shows 3 qubits suffering from $\sigma_{X}$ error labelled blue. Then, there will be 2 failures in the vertex stabilizer $A^{1}_{\alpha}$ checks, where the corresponding vertices are labelled red. Figure 8(b) shows that 2 qubits suffer from $\sigma_{Z}$ error labelled blue, and in this case, there will be 6 failures in the face stabilizers, $B^{1}_{\alpha}$, where the corresponding faces are labelled red. Notice that for both cases, the stabilizer failures always appear in an even number of sites; thus, the error correction process can be theoretically done via different matching functions, as discussed above [40, 11].
Figure 7 illustrates the logical $\sigma_{X}$ and $\sigma_{Z}$ graphically. The blue lines in Figure 7 forms a non-contractible loop $\gamma^{1}\in C_{1}(T^{3};\mathbb{Z}_{2}),\partial\gamma^{1}=0$ and $\gamma^{1}]\neq 0$. The red lines forms a non-compressible 2-torus $\gamma^{2}\in C_{2}(T^{3};\mathbb{Z}_{2}),\partial\gamma^{2}=0$ and $\gamma^{2}]\neq 0$. And they intersect at exactly 1 site, which is the edge drawn in green. These properties allow us to represent a logical $\sigma_{X}$ on the blue qubits and a logical $\sigma_{Z}$ on the red qubits.
@@ IV.3 Implementing 3-Torus Code: Encoding on Faces
Figure 9: The stabilizer checks defined on the 3-torus code for 2-dimensional cells. For each edge, we check the $\sigma_{Z}$ of 4 qubits reside on adjacent faces, the $A_{i}$ stabilizers, drawn in blue. For each cube, we check the $\sigma_{X}$ of 6 qubits residing on the boundary faces, the $B_{j}$ stabilizers, drawn in red.
@@
We have shown above that the 3-torus code can encode 3 qubits on the 1-cells. Here, we demonstrate that the 3-torus code can encode 3 extra qubits on the 2-cells, or faces. Note that the 3 qubits encoded on the faces are independent of the 3 qubits that are on the edges.
@@
The edge stabilizer $A^{2}_{i}$ is defined to check the $\sigma_{Z}$ of the $4$ qubits on the faces adjacent to a given edge, an example is drawn in blue in Figure 9. The cube stabilizer $B^{2}_{j}$ is defined to check the $\sigma_{X}$ of the $6$ qubits on the boundary faces of a given cube; an example is drawn in red in Figure 9. $A_{i}^{2}$ and $B_{j}^{2}$ commute for all $i,j$, because they either do not have any qubits in common or have 2 qubits in common. Not surprisingly,
@@
Here, we consider the case of qubits encoded on edges. In a general $u\times v\times w$ lattice for $u,v,w\in\mathbb{Z}^{+}$, there are in total $4uvw$ edges in the lattice. Hence, there will be $4uvw$ physical qubits and three logical qubits can be encoded, where the $\sigma_{Z}$ distance of the 3 qubits equals $u,v$, and $w$ respectively. Furthermore, the $\sigma_{X}$ distance of the 3 qubits equals $vw,wu$, and $uv$ respectively, where Figure 7 has demonstrated the idea. In the special case of a cubic lattice with dimensions $N\times N\times N$, there are $4N^{3}$ edges, the distance for logical $\sigma_{Z}$ is $N$, and the distance for logical $\sigma_{X}$ is $N^{2}$. Hence, the code is a $4N^{3},3,N]]$.